Please help me solve these problems
FIND THE PRODUCT (6x-7)^2
FIND THE PRODUCT (8x-3z)(8x+3z)
USING THE FORMULA (a+b)(a-b)=a^2-b^2 to multiply (9x-7)(9x+7)

Answer by ewatrrr(24785)


Hi,
FIND THE PRODUCT (6x-7)^2
(6x-7)(6x-7)= 36x^2-84x + 49 (sum of inner product(-42x) and outer product(-42x)=-84x
FIND THE PRODUCT
(8x-3z)(8x+3z)= 64x^2 -9z^2 (sum of inner product(-24x) and outer product(24x) = 0
USING THE FORMULA (a+b)(a-b)=a^2-b^2 to multiply
(9x-7)(9x+7) = 81x^2 - 49 (sum of inner product(-63x) and outer product(63x) = 0
Note: If using the outer and inner products for the x term is too confusing,
we can alwyas resort to the 'long way'
(6x-7)(6x-7)= 6x(6x-7)-7(6x-7) = 36x^2 -42x -42x + 49 = 36x^2 -84x -49